2. (5 pts) Let f(x)= and x+1 5 g(x)= x-3 Find and simplify the composite function f(g(x)). Write as a single fraction.
Answer #1

5 Sol) X-3 Given, f(x) = 2+18(x) = Now, (fog)(x) = f(g(x)) Flat 5 5 =f X- X-3 (: 8(x) = (:: f(x)=x71 1 = 5 +1 X-3 1 5 + x - 3